# Generative AI for Cel-Animation: A Survey

### v4

@@ 1 Introduction
Recent breakthroughs in generative artificial intelligence (GenAI), including large language models (LLMs) [118, 134, 56], multimodal LLMs (MLLMs) [68, 1, 110, 121, 17, 155, 41, 40, 142], and diffusion models [38, 99, 146, 29, 6], offer promising solutions to these challenges. By automating repetitive and technically demanding tasks, GenAI frees animators to engage more deeply in high-level creative processes. To contextualize how GenAI reshapes animation workflows, we review Cel-Animation’s evolution through three historical phases, as illustrated in Figure 1:
@@
The introduction of celluloid sheets enabled animators to separate moving characters from static backgrounds, significantly enhancing scene complexity and allowing efficient reuse of backgrounds [105]. Core animation techniques developed during this era include storyboarding, keyframe drawing, manual inbetweening, and hand-painting of individual cels. Landmark works, such as Disney’s Snow White and the Seven Dwarfs (1937), exemplify both the artistic potential and the inherent limitations of this approach: the film required over 200K hand-drawn frames and extensive human labor, highlighting issues of scalability and consistency.
@@
Advancements in digital technology significantly altered traditional Cel-Animation pipelines from the 1980s onward. Notably, Disney’s Computer Animation Production System (CAPS) [126], developed jointly with Pixar, automated coloring and compositing, achieving higher precision and complexity exemplified by films such as Beauty and the Beast (1991) and The Lion King (1994). Concurrently, Japanese studios leveraged tools such as Toonz [117], OpenToonz [82], Adobe Flash [125], and Clip Studio Paint [11] to streamline inbetweening and compositing, prominently demonstrated by productions like Nausicaä of the Valley of the Wind (1984) and Neon Genesis Evangelion (1995). Despite these innovations, digital technologies primarily served as supportive tools rather than automated solutions, still heavily relying on manual artistic inputs.
@@
Unlike prior technological innovations, Generative AI actively participates in the animation pipeline, significantly automating tasks traditionally requiring intensive human effort. Recent AI-driven methods automate complex processes, including inbetween frame generation, coloring, complex visual effects, and storyboard production [34, 77, 132]. For instance, AniDoc [77] employs video diffusion models to automate frame interpolation and colorization, and ToonCrafter [132] handles exaggerated non-linear movements and occlusions characteristic of cartoon animations. Platforms like Storyboarder.ai facilitate rapid pre-production by generating detailed storyboards directly from text prompts [104]. Practical implementations, such as Netflix Japan’s experimental short film The Dog and the Boy (2023), demonstrate the viability of AI-driven workflows in significantly reducing resource requirements. Furthermore, advanced video-generation frameworks like AniSora [49] underline the transformative potential of AIGC in enhancing both artistic creativity and production scalability.
Despite significant progress, AIGC integration in animation continues to face critical challenges, including maintaining stylistic consistency, ensuring temporal coherence, and precisely capturing human creative intent [34, 77, 132, 101]. Addressing these challenges requires systematic examination and thoughtful integration of GenAI tools within established animation practices.
Several recent surveys have explored related areas in generative media production. Li et al. [59] discuss challenges in long-video generation, while Lei et al. [58] and Xing et al. [133] examine realistic human motion synthesis and video diffusion models, respectively. Zhou et al. [154] focus on the broader applications of generative AI in video creation, and Cho et al. [19] extensively review text-to-video generation methods. Zhao et al. [152] offer insights into static cartoon image processing with limited exploration of animated contexts. Zhang et al. [147] examine the recent advances in GenAI for film creation. In contrast, our survey explicitly addresses GenAI’s integration across the complete Cel-Animation pipeline, examining its impacts on pre-production, production, and post-production phases, alongside critical discussions of existing challenges and promising research directions.
@@ 2.2.1 Large Language Models (LLMs)
  • Scaling Laws [52]: Model performance grows predictably with increased parameters, data, and compute.
  • Emergent Abilities [151]: At scale, models exhibit novel behaviors like in-context learning, instruction following, and chain-of-thought reasoning.
LLMs have been applied in animation and filmmaking to streamline tasks like scriptwriting, character backstory creation, and generating visual descriptions for storyboarding. MLLMs [155, 128] extend LLMs by integrating visual encoders and cross-modal aligners, excelling in tasks that require both visual and textual reasoning.
@@ 2.2.2 GAN, VAE, and Diffusion Models
Generative Adversarial Networks (GANs) [33] consist of generators and discriminators trained adversarially, applicable to background creation and style transfer (StyleGAN [54], Pix2Pix [47]). Variational Autoencoders (VAEs) [55] optimize latent variable distributions, effective for generating controllable poses and expressions. Diffusion models [98, 38] progressively denoise samples via:
@@
predicting noise with neural networks $\epsilon_{\theta}(x_{t},t)$. Extensions like Stable Diffusion [87] and ControlNet [146] support animation-specific tasks, addressing temporal consistency.
@@
Pre-production 3.1) Scripting (§3.1) HoLLMwood [14], ChatGPT [1], Gemini [86], Claude [4], LLaMA [4], InternVL [18], Qwen [5] Setting (§3.1) Midjourney [79], Stable Diffusion [87] Storyboarding (§3.1) StoryWeaver [143], CogCartoon [157], Coherent Story [84], Seed-Story [138], Make-A-Story [85], Make-A-Storyboard [107], Storyboarder.ai [104], Animate-A-Story [37], StoryImager [112], VideoDirectorGPT [64], CustomCrafter [130], VideoStudio [73], Mind the Time [131], DREAMRUNNER [123], Vlogger [158], Anim-Director [62], StoryGen [66], Production 3.2) Layout (L/O) (§3.2) CogCartoon [157], SGSIG [148], VideoComposer [122], LayoutGAN [60], Layout2Vid [64], DiffSensei [129], MangaDiffusion [16], CameraCtrl [36] Keyframe Animation (§3.2) AnimateAnyone [39], Champ [156], MimicMotion [149], Animate-X [108], MikuDance [144], CoNR [65] Inbetweening (§3.2) ToonCraft [132], JointStroke [80], AniSora [49], AnimeInterp [96], AnimeInbet [30], AutoFI [92], SAIN [91], MotionBridge [111], toona.io [116], DeepSketch-Guided [61], DeepGeometrized [97], LayerAnimate [140] Colorization (§3.2) AniDoc [78], ToonCraft [132], VToonify [135], TokenFlow [32], StyleGANEX [136], FRESCO [139], toona.io [116], TRE-Net [120], LayerAnimate [140], DeepLineArt [95], SketchBetween [72], Flow Estimate-Refine Net [141], PromptFix [142], MangaNinja [71] , Animation Xformer [10], Inclusion Matching [28], LVCD [45] Post-production 3.3) Compositing & Photography (§3.3) IC-Light [3], Dovenet [24], High res harmonization [25], Pct-net [35], SSH [48], Object Compositing [113], SSN [93], Dr.Bokeh [94], Floating No More [75], ObjectDrop [127], Alchemist [88], Disenstudio [13], LayerAnimate [140] Cutting (CT) (§3.3) M-SAN [109], Reframe Anything [8], OpusClip [83] Music & Sound Effects (§3.3) FoleyMusic [31], Video2Music [51], V2Meow [106], MelFusion [20], VidMuse [114], SpecVQGAN [46], Im2Wav [89], FoleyGen [76], DiffFoley [74], Action2Sound [12] After Record (AR) & Dubbing (DB) (§3.3) StyleDubber [23], ANIM-400K [7], EmoDubber [22], MovieDubbing [150], HPMDubbing [21], V2C [15] Others (§3.4) Cel-Animation Editing (§3.4) Re:Draw [9], ScalingConcept [42] Cels Decomposition 3.4) Sprite-from-Sprite [145], Generative Omnimatte [57], TransPixar [119], LayerAnimate [140] 3D Assistance (§3.4) Toon3D [124], CoNR [65], DrawingSpinUp [153]
@@ 3.1 GenAI for Pre-production
LLMs facilitate the creation of story settings, plot expansion, and dialogue generation for original animations. For adaptations (e.g., novels, manga, games), LLMs first comprehend source content including narrative structures and character dynamics before adapting or generating scripts. Recent proprietary (ChatGPT [1], Claude [4], Gemini [86]) and open-source LLMs (LLaMA [118], Qwen [5]) demonstrate substantial adaptability in both original and adapted animation scripts. Frameworks like HoLLMwood [14] automate scriptwriting processes, enhancing narrative coherence and character interaction quality.
@@
Dedicated research for automated setting generation is limited, but practical use cases demonstrate tools such as Midjourney [79] and Stable Diffusion [87] effectively generating characters, scenes, and objects from text prompts. As shown in Figure 4, these generated contents demonstrate a certain degree of consistency (for example, multiple views of characters from different angles maintain good identity preservation). However, detailed inaccuracies remain common, including irrelevant items or nonsensical text within generated designs.
Figure 4: Examples of AI-generated Settings with Midjourney [79] and Stable Diffusion [87], showing various designs including items, characters, scenes, and mecha.
@@
Recent GenAI techniques address automated storyboard generation primarily for live-action films [64] and non-photorealistic animations [157, 143]. Typically, textual scripts are processed by LLMs into detailed scene descriptions, which then guide image generation models to produce sequential visual narratives. Core challenges include maintaining consistent character identity and ensuring alignment between human intent and AI-generated content. Strategies addressing these challenges involve memory-based identity retention [85, 84], attention-based controls [138], and semantic alignment enhancements [90, 143]. CogCartoon [157] further improves storyboard precision by explicitly controlling character positions within frames.
@@ 3.2 GenAI for Production
L/O phase refines rough storyboards by defining character positions, scene composition, perspective, camera angles, and lighting. While some storyboard methods support bounding-box-based spatial control, they often lack fine-grained manipulation of poses and orientations [60]. Recent approaches [148] improve layout realism by taking sketches as input, enabling control over object size, pose, depth, and perspective. However, these models primarily target object layouts in photorealistic images, limiting effectiveness for anime-specific character layouts. Moreover, layout generation remains largely static. CameraCtrl [36] introduces camera motion control in video generation, enabling dynamic layout synthesis with synchronized foreground-background motion and temporal continuity.
@@
Keyframe animation in Cel-Animation focuses on animating scene elements, primarily characters. Recent methods [39, 53] use 2D body landmarks to guide pose generation with spatial fidelity. Systems like Animate-Any-One [39], Champ [156], and MusePose [115] introduce ReferenceNet to bridge source-target pose differences, enhancing transfer accuracy and realism. MimicMotion [149] and Animate-X [108] further generalize across body types, including cartoon characters, by modeling morphological variance. For facial animation, recent works [101, 102, 100, 70] animate faces using 3D Gaussians or neural representations. TextToon [101] and Emo-Avatar [70] integrate LLMs with 3D Gaussian splatting for controllable cartoon stylization. Ada-TalkingHead [102] uses neural keypoints for head animation, while Editable-Head [100] adopts landmark-based approaches. These models offer greater control and finer details in facial motion compared to coarse full-body animation methods.
@@
While mechanically straightforward, inbetweening is among the most labor-intensive and time-consuming tasks in animation, making it ideal for GenAI acceleration. Accordingly, significant research has emerged in this area. State-of-the-art approaches like ToonCraft [132] and AutoFI [92] adapt video interpolation techniques using diffusion models to generate intermediate frames between keyframes. Other methods incorporate user guidance or structural features: [61] uses sketches to preserve motion and visual style; SAIN [91] employs multi-stream U-Transformers to capture region, stroke, and pixel-level cues; AnimeInbet [97] models line drawings as vector graphs for structure-aware inbetweening. Interactive approaches such as [30] provide trajectory-guided sliders for fine control, while [80] aligns stroke structures across frames. MotionBridge [111] further enhances controllability and temporal coherence using structured temporal-spatial representations.
@@
Colorization is another highly manual process extensively studied in the anime domain. Our focus is on video-level methods using line drawings and sparse color references. Classical models [77] leverage a single colored drawing to guide colorization of black-and-white inputs. Newer methods treat this as a style transfer task, with few-shot approaches like TextToon [101], StyleGANEX [136], and VToonify [135] learning sketch-to-style mappings. However, these often struggle with temporal consistency and generalize poorly beyond avatars. Recent works [137, 139, 32] use large pretrained diffusion models to achieve zero-shot, temporally consistent video stylization. Additionally, scene-level control has improved with models [10, 45, 141, 28] offering detailed and controllable colorization. Some systems, such as ToonCraft [132], support simultaneous inbetweening and colorization using only sparse input (e.g., one colored frame plus keyframes).
Figure 5: Comparison of colorization generated by GenAI models. The example is adapted from [78].

### v5

@@ 1 Introduction
Recent breakthroughs in generative artificial intelligence (GenAI), including large language models (LLMs) [touvron2023llama, yang2024qwen2, le2023bloom], multimodal LLMs (MLLMs) [liu2024visual, achiam2023gpt, tang2023video, Qwen2VL, chen2024far, zhu2023minigpt, hua2024mmcomposition, hua2024finecaption, yu2024promptfix], and diffusion models [ho2020ddpm, song2020denoising, zhang2023controlnet, esser2024scaling, blattmann2023align], offer promising solutions to these challenges. By automating repetitive and technically demanding tasks, GenAI frees animators to engage more deeply in high-level creative processes. To contextualize how GenAI reshapes animation workflows, we review Cel-Animation’s evolution through three historical phases, as illustrated in Figure 1:
@@
The introduction of celluloid sheets enabled animators to separate moving characters from static backgrounds, significantly enhancing scene complexity and allowing efficient reuse of backgrounds [studiobinder_cel_animation]. Core animation techniques developed during this era include storyboarding, keyframe drawing, manual inbetweening, and hand-painting of individual cels. Landmark works, such as Disney’s Snow White and the Seven Dwarfs (1937), exemplify both the artistic potential and the inherent limitations of this approach: the film required over 200K hand-drawn frames and extensive human labor, highlighting issues of scalability and consistency.
@@
Advancements in digital technology significantly altered traditional Cel-Animation pipelines from the 1980s onward. Notably, Disney’s Computer Animation Production System (CAPS) [caps_wikipedia], developed jointly with Pixar, automated coloring and compositing, achieving higher precision and complexity exemplified by films such as Beauty and the Beast (1991) and The Lion King (1994). Concurrently, Japanese studios leveraged tools such as Toonz [toonz_wikipedia], OpenToonz [opentoonz], Adobe Flash [adobeflash], and Clip Studio Paint [clipstudio] to streamline inbetweening and compositing, prominently demonstrated by productions like Nausicaä of the Valley of the Wind (1984) and Neon Genesis Evangelion (1995). Despite these innovations, digital technologies primarily served as supportive tools rather than automated solutions, still heavily relying on manual artistic inputs.
@@
Unlike prior technological innovations, Generative AI actively participates in the animation pipeline, significantly automating tasks traditionally requiring intensive human effort. Recent AI-driven methods automate complex processes, including inbetween frame generation, coloring, complex visual effects, and storyboard production [guajardo2024generativeai, anidoc2024, tooncrafter2024]. For instance, AniDoc [anidoc2024] employs video diffusion models to automate frame interpolation and colorization, and ToonCrafter [tooncrafter2024] handles exaggerated non-linear movements and occlusions characteristic of cartoon animations. Platforms like Storyboarder.ai facilitate rapid pre-production by generating detailed storyboards directly from text prompts [storyboarder]. Practical implementations, such as Netflix Japan’s experimental short film The Dog and the Boy (2023), demonstrate the viability of AI-driven workflows in significantly reducing resource requirements. Furthermore, advanced video-generation frameworks like AniSora [jiang2024anisora] underline the transformative potential of AIGC in enhancing both artistic creativity and production scalability.
Despite significant progress, AIGC integration in animation continues to face critical challenges, including maintaining stylistic consistency, ensuring temporal coherence, and precisely capturing human creative intent [guajardo2024generativeai, anidoc2024, tooncrafter2024, song2024texttoon]. Addressing these challenges requires systematic examination and thoughtful integration of GenAI tools within established animation practices.
Several recent surveys have explored related areas in generative media production. Li et al. [li2024survey] discuss challenges in long-video generation, while Lei et al. [lei2024survey] and Xing et al. [xing2024survey] examine realistic human motion synthesis and video diffusion models, respectively. Zhou et al. [zhou2024survey] focus on the broader applications of generative AI in video creation, and Cho et al. [cho2024survey] extensively review text-to-video generation methods. Zhao et al. [Zhao2022Cartoon] offer insights into static cartoon image processing with limited exploration of animated contexts. Zhang et al. [zhang2025generative] examine the recent advances in GenAI for film creation. In contrast, our survey explicitly addresses GenAI’s integration across the complete Cel-Animation pipeline, examining its impacts on pre-production, production, and post-production phases, alongside critical discussions of existing challenges and promising research directions.
@@ 2.2.1 Large Language Models (LLMs)
  • Scaling Laws [kaplan2020scaling]: Model performance grows predictably with increased parameters, data, and compute.
  • Emergent Abilities [zhao2023survey]: At scale, models exhibit novel behaviors like in-context learning, instruction following, and chain-of-thought reasoning.
LLMs have been applied in animation and filmmaking to streamline tasks like scriptwriting, character backstory creation, and generating visual descriptions for storyboarding. MLLMs [zhu2023minigpt, liu2023visual] extend LLMs by integrating visual encoders and cross-modal aligners, excelling in tasks that require both visual and textual reasoning.
@@ 2.2.2 GAN, VAE, and Diffusion Models
Generative Adversarial Networks (GANs) [goodfellow2014generative] consist of generators and discriminators trained adversarially, applicable to background creation and style transfer (StyleGAN [karras2019style], Pix2Pix [isola2017image]). Variational Autoencoders (VAEs) [kingma2013auto] optimize latent variable distributions, effective for generating controllable poses and expressions. Diffusion models [sohl2015deep, ho2020ddpm] progressively denoise samples via:
@@
predicting noise with neural networks $\epsilon_{\theta}(x_{t},t)$. Extensions like Stable Diffusion [rombach2022high] and ControlNet [zhang2023controlnet] support animation-specific tasks, addressing temporal consistency.
@@
Pre-production 3.1) Scripting (§3.1) HoLLMwood [chen2024hollmwood], ChatGPT [achiam2023gpt], Gemini [reid2024gemini], Claude [anthropic_claude35], LLaMA [anthropic_claude35], InternVL [chen2024internvl], Qwen [bai2023qwen] Setting (§3.1) Midjourney [midjourney], Stable Diffusion [rombach2022high] Storyboarding (§3.1) StoryWeaver [zhang2024storyweaver], CogCartoon [cogcartoon2023], Coherent Story [Pan_2024_WACV], Seed-Story [yang2024seed], Make-A-Story [rahman2023make], Make-A-Storyboard [su2023make], Storyboarder.ai [storyboarder], Animate-A-Story [he2023animate-a-story], StoryImager [tao2025storyimager], VideoDirectorGPT [lin2023videodirectorgpt], CustomCrafter [wu2024customcrafter], VideoStudio [long2024videodrafter], Mind the Time [wu2024mind], DREAMRUNNER [wang2024dreamrunner], Vlogger [zhuang2024vlogger], Anim-Director [li2024anim], StoryGen [liu2024storygen], Production 3.2) Layout (L/O) (§3.2) CogCartoon [cogcartoon2023], SGSIG [zhang2024sketch], VideoComposer [wang2024videocomposer], LayoutGAN [li2019layoutgan], Layout2Vid [lin2023videodirectorgpt], DiffSensei [wu2024diffsensei], MangaDiffusion [chen2024manga], CameraCtrl [he2024cameractrl] Keyframe Animation (§3.2) AnimateAnyone [hu2023animateanyone], Champ [zhu2024champ], MimicMotion [mimicmotion2024], Animate-X [tan2024animate], MikuDance [zhang2024mikudance], CoNR [lin2022collaborative] Inbetweening (§3.2) ToonCraft [tooncrafter2024], JointStroke [JointStroke], AniSora [jiang2024anisora], AnimeInterp [siyao2021deep], AnimeInbet [fukusato2024exploring], AutoFI [shen2022autofi], SAIN [shen2024bridging], MotionBridge [tanveer2024motionbridge], toona.io [toonaio], DeepSketch-Guided [li2021deep], DeepGeometrized [siyao2023deep], LayerAnimate [yang2025layeranimate] Colorization (§3.2) AniDoc [meng2024anidoc], ToonCraft [tooncrafter2024], VToonify [yang2022Vtoonify], TokenFlow [tokenflow2023], StyleGANEX [yang2023styleganex], FRESCO [yang2024fresco], toona.io [toonaio], TRE-Net [TRE-Net], LayerAnimate [yang2025layeranimate], DeepLineArt [shi2020deeplineart], SketchBetween [loftsdottir2022sketchbetween], Flow Estimate-Refine Net [Yu2022AnimationColorization], PromptFix [yu2024promptfix], MangaNinja [liu2025manganinja] , Animation Xformer [casey2021animation], Inclusion Matching [dai2024learning], LVCD [huang2024lvcd] Post-production 3.3) Compositing & Photography (§3.3) IC-Light [zhang2024iclight], Dovenet [cong2020dovenet], High res harmonization [cong2022high], Pct-net [guerreiro2023pct], SSH [jiang2021ssh], Object Compositing [tarres2024thinking], SSN [sheng2021ssn], Dr.Bokeh [sheng2024dr], Floating No More [man2024floating], ObjectDrop [winter2024objectdrop], Alchemist [sharma2024alchemist], Disenstudio [chen2024disenstudio], LayerAnimate [yang2025layeranimate] Cutting (CT) (§3.3) M-SAN [Tang_2022_ACCV], Reframe Anything [cao2024reframe], OpusClip [opusclip] Music & Sound Effects (§3.3) FoleyMusic [FoleyMusic2020], Video2Music [KANG2024123640], V2Meow [su2024v2meow], MelFusion [chowdhury2023melfusion], VidMuse [tian2024vidmuse], SpecVQGAN [SpecVQGAN_Iashin_2021], Im2Wav [sheffer2022i], FoleyGen [mei2024foleygen], DiffFoley [luo2024diff], Action2Sound [chen2024action2sound] After Record (AR) & Dubbing (DB) (§3.3) StyleDubber [cong2024styledubber], ANIM-400K [cai2024anim400k], EmoDubber [cong2024emodubberhighqualityemotion], MovieDubbing [zhang2024from], HPMDubbing [cong2023learning], V2C [chen2022v2c] Others (§3.4) Cel-Animation Editing (§3.4) Re:Draw [cardoso2024re], ScalingConcept [huang2024scaling] Cels Decomposition 3.4) Sprite-from-Sprite [zhang2022sprite], Generative Omnimatte [lee2024generativeOmnimatte], TransPixar [wang2025transpixar], LayerAnimate [yang2025layeranimate] 3D Assistance (§3.4) Toon3D [weber2024toon3d], CoNR [lin2022collaborative], DrawingSpinUp [zhou2024drawingspinup]
@@ 3.1 GenAI for Pre-production
LLMs facilitate the creation of story settings, plot expansion, and dialogue generation for original animations. For adaptations (e.g., novels, manga, games), LLMs first comprehend source content including narrative structures and character dynamics before adapting or generating scripts. Recent proprietary (ChatGPT [achiam2023gpt], Claude [anthropic_claude35], Gemini [reid2024gemini]) and open-source LLMs (LLaMA [touvron2023llama], Qwen [bai2023qwen]) demonstrate substantial adaptability in both original and adapted animation scripts. Frameworks like HoLLMwood [chen2024hollmwood] automate scriptwriting processes, enhancing narrative coherence and character interaction quality.
@@
Dedicated research for automated setting generation is limited, but practical use cases demonstrate tools such as Midjourney [midjourney] and Stable Diffusion [rombach2022high] effectively generating characters, scenes, and objects from text prompts. As shown in Figure 4, these generated contents demonstrate a certain degree of consistency (for example, multiple views of characters from different angles maintain good identity preservation). However, detailed inaccuracies remain common, including irrelevant items or nonsensical text within generated designs.
Figure 4: Examples of AI-generated Settings with Midjourney [midjourney] and Stable Diffusion [rombach2022high], showing various designs including items, characters, scenes, and mecha.
@@
Recent GenAI techniques address automated storyboard generation primarily for live-action films [lin2023videodirectorgpt] and non-photorealistic animations [cogcartoon2023, zhang2024storyweaver]. Typically, textual scripts are processed by LLMs into detailed scene descriptions, which then guide image generation models to produce sequential visual narratives. Core challenges include maintaining consistent character identity and ensuring alignment between human intent and AI-generated content. Strategies addressing these challenges involve memory-based identity retention [rahman2023make, Pan_2024_WACV], attention-based controls [yang2024seed], and semantic alignment enhancements [shen2024boosting, zhang2024storyweaver]. CogCartoon [cogcartoon2023] further improves storyboard precision by explicitly controlling character positions within frames.
@@ 3.2 GenAI for Production
L/O phase refines rough storyboards by defining character positions, scene composition, perspective, camera angles, and lighting. While some storyboard methods support bounding-box-based spatial control, they often lack fine-grained manipulation of poses and orientations [li2019layoutgan]. Recent approaches [zhang2024sketch] improve layout realism by taking sketches as input, enabling control over object size, pose, depth, and perspective. However, these models primarily target object layouts in photorealistic images, limiting effectiveness for anime-specific character layouts. Moreover, layout generation remains largely static. CameraCtrl [he2024cameractrl] introduces camera motion control in video generation, enabling dynamic layout synthesis with synchronized foreground-background motion and temporal continuity.
@@
Keyframe animation in Cel-Animation focuses on animating scene elements, primarily characters. Recent methods [hu2023animateanyone, karras2023dreampose] use 2D body landmarks to guide pose generation with spatial fidelity. Systems like Animate-Any-One [hu2023animateanyone], Champ [zhu2024champ], and MusePose [musepose] introduce ReferenceNet to bridge source-target pose differences, enhancing transfer accuracy and realism. MimicMotion [mimicmotion2024] and Animate-X [tan2024animate] further generalize across body types, including cartoon characters, by modeling morphological variance. For facial animation, recent works [song2024texttoon, song2024adaptive, song2021talking, liu2024emo] animate faces using 3D Gaussians or neural representations. TextToon [song2024texttoon] and Emo-Avatar [liu2024emo] integrate LLMs with 3D Gaussian splatting for controllable cartoon stylization. Ada-TalkingHead [song2024adaptive] uses neural keypoints for head animation, while Editable-Head [song2021talking] adopts landmark-based approaches. These models offer greater control and finer details in facial motion compared to coarse full-body animation methods.
@@
While mechanically straightforward, inbetweening is among the most labor-intensive and time-consuming tasks in animation, making it ideal for GenAI acceleration. Accordingly, significant research has emerged in this area. State-of-the-art approaches like ToonCraft [tooncrafter2024] and AutoFI [shen2022autofi] adapt video interpolation techniques using diffusion models to generate intermediate frames between keyframes. Other methods incorporate user guidance or structural features: [li2021deep] uses sketches to preserve motion and visual style; SAIN [shen2024bridging] employs multi-stream U-Transformers to capture region, stroke, and pixel-level cues; AnimeInbet [siyao2023deep] models line drawings as vector graphs for structure-aware inbetweening. Interactive approaches such as [fukusato2024exploring] provide trajectory-guided sliders for fine control, while [JointStroke] aligns stroke structures across frames. MotionBridge [tanveer2024motionbridge] further enhances controllability and temporal coherence using structured temporal-spatial representations.
@@
Colorization is another highly manual process extensively studied in the anime domain. Our focus is on video-level methods using line drawings and sparse color references. Classical models [anidoc2024] leverage a single colored drawing to guide colorization of black-and-white inputs. Newer methods treat this as a style transfer task, with few-shot approaches like TextToon [song2024texttoon], StyleGANEX [yang2023styleganex], and VToonify [yang2022Vtoonify] learning sketch-to-style mappings. However, these often struggle with temporal consistency and generalize poorly beyond avatars. Recent works [yang2023rerender, yang2024fresco, tokenflow2023] use large pretrained diffusion models to achieve zero-shot, temporally consistent video stylization. Additionally, scene-level control has improved with models [casey2021animation, huang2024lvcd, Yu2022AnimationColorization, dai2024learning] offering detailed and controllable colorization. Some systems, such as ToonCraft [tooncrafter2024], support simultaneous inbetweening and colorization using only sparse input (e.g., one colored frame plus keyframes).
Figure 5: Comparison of colorization generated by GenAI models. The example is adapted from [meng2024anidoc].